# Heuristic search for linear positive systems

### v1

@@ Abstract
This work considers infinite-horizon optimal control of positive linear systems applied to the case of network routing problems. We demonstrate the equivalence between Stochastic Shortest Path (SSP) problems and optimal control of a certain class of linear systems. This is used to construct a heuristic search framework for this class of linear systems inspired by existing methods for SSP. We propose a heuristics-based algorithm for finding local solutions to the analyzed class of optimal control problems with positive state and linear dynamics. More fundamentally, the results allow for analysis of the conditions for explicit solutions to the Bellman equation utilized by heuristic search methods.
@@ 1 Introduction
The work presented here demonstrates the equivalence of two problem classes that are the focus of optimal control and graph search algorithms respectively. The first is a class of positive linear systems with linear cost and homogeneous constraints. The second is the class of Stochastic Shortest Path (SSP) problems and their generalizations, which are the focus of contemporary work on heuristic search algorithms, see (Bonet & Hansen, 2010) for a summary. In bridging the gap between these two formulations, we show that many results derived for either can be readily applied to the other. Specifically, we give a framework for a heuristics-based solution of optimal control problems with continuous state and action space that belong to the above described class of positive systems. In a similar spirit to heuristic search algorithms developed for SSP, these methods are computationally feasible for high-dimensional state and action spaces. The analogy between the two classes also gives an analytic framework for further development of existing methods for SSP using results within the theory of optimal control.
@@
The method of Dynamic Programming (DP), developed in (Bellman, 1957), shares a common origin with Dijkstra’s algorithm, see (Dijkstra, 1959). As emphasized in (Sniedovich, 2006), the line of argument in Dijkstra’s original work closely mirrors the optimality principle as first formulated by Bellman. Since their inception, these methods have both gone on to inspire a wide range of extensions, seeing great success in their respective fields of control theory and computer science. Although the lines of work within each discipline may seem disparate, they share the common feature of utilizing explicit solutions to the Bellman equation to achieve computational efficiency. It is this feature that enables the present work to give favorable results for positive systems and interpret these in terms of their counterpart in the SSP formulation.
@@
A natural extension to Dijkstra’s algorithm is proposed in (Hart, Nilsson & Raphael, 1968). The authors introduce a heuristic function that lower bounds the cost in any state and uses this to guide the search, prioritizing directions that give a low heuristic cost. Already in (Bertsekas & Tsitsiklis, 1991) the theoretical results of DP obtained in the deterministic case are extended to stochastic problems. Algorithms that implement similar ideas beyond the deterministic case through the use of heuristics are the subject of subsequent work in the field, see the seminal paper (Barto, Bradtke & Singh, 1995). The close underlying connection between the study of heuristic functions and Lyapunov analysis is highlighted in (Perkins & Barto, 2001). Comprehensive theoretical results on control of the broader class of constrained Markov Decision Processes (MDP), which includes problems like SSP, can be found in the monograph (Altman, 1999). Therein, arguments based in DP and Lyapunov theory are used to formulate linear programs for problems in this class under certain conditions.
@@
In (McMahan, Likhachev & Gordon, 2005) the authors incorporate the use of a heuristic upper bound and provide methods to find and improve such a bound. The extension is motivated by computational advantages, since the total cost is known to be upper bounded from the problem formulation of SSP. In this paper, the upper bound will additionally serve the purpose of guaranteeing the stability of solutions, which is not otherwise certain in the setting of optimal control. Experimental results demonstrating the robust performance of heuristic search methods for problems that deviate from the prescribed structure can be found in (McMahan & Gordon, 2005), together with an extended discussion of priority calculations for this family of algorithms. (Smith & Simmons, 2006) seek to improve existing algorithms by using upper and lower bound heuristics to inform the search. The class of problems for which heuristic methods can be applied is expanded in (Kolobov, Mausam, Weld & Geffner, 2011) by the introduction of a broader class of SSP and a corresponding algorithm. The recent work (Schmoll & Schubert, 2018) further improves existing methods for certain classes of problems that can be rewritten as SSP.
@@
Within the field of optimal control, recent works like (Blanchini, Bolzern, Colaneri, De Nicolao & Giordano, 2023) and (Rantzer, 2022) focus on explicit solutions to the Bellman equation for subclasses of positive linear systems. In (Ohlin, Tegling & Rantzer, 2024), the class of applicable systems is extended to include coupled homogeneous input constraints, allowing for the modeling of network routing problems while maintaining these computational advantages. Similar properties of solutions to the Bellman equation for a set of MDP with cost function based on the Kullback-Leibler divergence are noted in (Todorov, 2006), with applications to optimal transport on networks. The use of upper and lower bounds on a general cost function for optimal control problems is explored in (Lincoln & Rantzer, 2006) as a means to relax the problem formulation, allowing solutions to be suboptimal to an extent, in exchange for improved performance.
@@ 1.1 Notation
[…]
@@ 2 Problem setup
with ${x_{0}\in\mathbb{R}^{n}_{+}}$. The matrices ${B=\begin{bmatrix}B_{1}\cdots B_{n}\end{bmatrix}\in\mathbb{R}^{n\times m}}$ and ${A\in\mathbb{R}^{n\times n}}$ with ${B_{i}\in\mathbb{R}^{n\times m_{i}}}$ define the linear dynamics. The input signal ${u\in\mathbb{R}^{m}}$ is partitioned into $n$ subvectors $u_{i}$, each containing $m_{i}$ elements, so that ${m=\sum_{i=1}^{n}m_{i}}$. Let the matrices $B_{i}$ be further subdivided into individual columns as ${B_{i}=\begin{bmatrix}B_{i1}\cdots B_{im_{i}}\end{bmatrix}}$ with ${B_{ij}\in\mathbb{R}^{n}}$. The costs connected to the states and actions are ${s\in\mathbb{R}^{n}_{+}}$ with ${s>0}$ and ${r\in\mathbb{R}^{m}_{+}}$ with ${r_{i}\in\mathbb{R}^{m_{i}}_{+}}$ following the partition of $u$ and $r_{ij}$ denoting the $j$th element of $r_{i}$. The constraints on the input signal $u(t)$ are given by the matrix ${E=\begin{bmatrix}E_{1}\cdots E_{n}\end{bmatrix}^{\top}\in\mathbb{R}^{n\times n}_{+}}$, which we assume to be invertible. The special case of singular $E$ typically corresponds to degenerate formulations and will not be treated here. Define the set $\mathcal{V}$ to contain the indices $\{1,\ldots,n\}$. Let ${K=\left[K_{1}^{\top}\cdots K_{n}^{\top}\right]^{\top}}$ be a feedback matrix with ${K_{i}\in\mathbb{R}^{m_{i}\times n}_{+}}$ and define the set
@@
The above formulation with strictly positive $s$ ensures that the system is detectable with the specified cost function, meaning that the optimal feedback law must also be stabilizing. A more intricate formulation of this criterion is given in (Li & Rantzer, 2023). In the present setting, however, any detectable positive system on the form (1) can be transformed to an equivalent problem satisfying $s>0$ by reallocation of the costs $s$ and $r$.
@@ 2.1 Positive networks with linear cost
for all $K\in\mathcal{K}$ and all reachable states $x\in\mathbb{R}^{n}_{+}$.
@@
((Ohlin et al., 2024)) Under Assumption 1, the following three statements are equivalent:
  • ($i$) The problem (1) has a finite value for $x_{0}\in\mathbb{R}^{n}_{+}$.
  • ($ii$) There exists $p\in\mathbb{R}^{n}_{+}$ satisfying the equation
@@
  • ($iii$) The value of the linear program
@@
The proof, which is a modified version that in (Ohlin et al., 2024), can be found in the Appendix.
@@
$p=s+A^{\top}p+\min_{K\in\mathcal{K}}\left[K^{\top}(r+B^{\top}p)\right]$
where the dependence on the constraint matrices $E_{i}$ is built into the definition of the set $\mathcal{K}$. The optimal cost is attained by choosing $K$ as in (3). Although less explicit, this expression is more compact will be used when convenient in the treatment of the heuristic search algorithm presented in Sections 4 and 5 to illustrate the use of different policies for subsets of the state space $\mathcal{V}$.
@@ 3 The equivalent problem
The extension of previous results on positive linear systems to the class of systems covered by Theorem 1 lets us make a formal connection to other existing frameworks for network routing problems. As the definitions used in the literature concerning SSP (e.g. (Bertsekas & Tsitsiklis, 1991), (McMahan et al., 2005) and (Kolobov et al., 2011)) differ in interpretation, the following definition strives to be as broad as possible while maintaining applicability of the results in Theorem 1.
@@ A Stochastic Shortest Path (SSP) problem is defined by the following properties:
    a cost function $\mathcal{C}(v,a)$ associating each action with a non-negative cost that satisfies $\mathcal{C}(v,a)=0$ for $v\in\mathcal{G}$ and $\mathcal{C}(v,a)>0$ for $v\in\mathcal{V}^{\prime}\setminus\mathcal{G}$
@@
Note that the requirement $\mathcal{C}(v,a)>0$ for $v\in\mathcal{V}^{\prime}\setminus\mathcal{G}$ is equivalent to assuming detectability in the setting (1). (Kolobov et al., 2011) lifts this restriction to allow zero-cost cycles in states outside $\mathcal{G}$ to be part of the solution. This corresponds to the optimal cost being attained by a non-stabilizing controller in the setting of (1). Such cases require special treatment and will not be considered in the present work.
@@ 3.1 Reformulation
The problem (1) can be recast as an SSP. In order to show this, we first introduce a helpful lemma and an assumption on the system dynamics. The following lemma simplifies the subsequent problem construction by allowing the choice $E=I$ without loss of generality.
@@
Consider a similarity transform $\hat{x}=Tx$ of the original state vector in (1). This gives rise to transformed constraints $\mathbf{1}^{\top}u_{i}\leq E_{i}^{\top}T^{-1}\hat{x}$. Enforcing $ET^{-1}=I$ yields $T=E$, which is a valid choice since $E$ is invertible. The original state $x$ is confined to +$\mathbb{R}^{n}_{+}$ under the dynamics of (1). Further, the matrix $E$ is nonnegative by construction. Hence, it holds that $\hat{x}\geq 0$ starting from $\hat{x}_{0}=Tx_{0}$. $\blacksquare$ Next, we introduce an assumption on substochasticity of the closed loop system in (1). As is shown below, both classes can be interpreted as SSP. Without this assumption, however, the problem can only be reformulated as an equivalent SSP given a relaxation of Definition 3.1, to allow for an infinite extension of the state space. The resulting problem can still be formulated as an optimization over finitely many actions, and the search methods introduced in subsequent sections, inspired by methods for SSP, are equally applicable regardless of this condition.
@@
The matrix ${(A+BK)}$ is column substochastic for all $K\in\mathcal{K}$.
@@
    The closed-loop system ${x(t+1)=(A+BK)x(t)}$ with $K\in\mathcal{K}$ is a transition from the state $x(t)$ encoded by the transition matrix $(A+BK)$. The resulting state $x(t+1)$ is interpreted as a probability distribution over the discrete states of $\mathcal{V}^{\prime}$, and we define the transition function $\mathcal{T}(v,a)$ as the $v$th column of $(A+BK)$ where $K$ is given by the choice of action $a$. For all states $v$ and actions $a$ such that ${||\mathcal{T}(v,a)||_{1}<1}$, a transition to $v_{g}$ with probability ${1-||\mathcal{T}(v,a)||_{1}}$ is added. All that remains for the construction to be complete is to define the action set ${\mathcal{A}(v_{g})=\{a_{g}\}}$, where $a_{g}$ gives unit transition probability to $v_{g}$.
@@ 3.2 Equivalence
Consider a problem on the form (1) with finite value in which the system satisfies Assumptions 1 and 2. It can be rewritten as an SSP according to Definition 3.1, with equal expected cost. The optimal solution to this SSP is also an optimal control law for the problem (1).
@@
By Lemma 3.2, it is sufficient to consider systems with $E=I$. Properties 1-6 of Definition 3.1 are directly fulfilled by the construction in Section 3.1. All that remains is to show that 7 also holds for the resulting SSP, that the expected cost is identical to that of the original problem and that the solution attains the optimal value of (1).
In order to show property 7 of Definition 3.1 we apply the method used in Chapter 2.3 of (Bertsekas & Tsitsiklis, 1996) to rewrite any discounted MDP as an SSP. In the system ${x(t+1)=(A+BK)x(t)}$, the quantity $||\mathcal{T}(v,a)||_{1}$ corresponds to the column sums of ${(A+BK)}$. The matrix ${(A+BK)}$ is Schur stable for some choices of ${K\in\mathcal{K}}$, since the problem (1) has a finite value. In particular, this is at least true for the optimal feedback law. By the Perron-Frobenius theorem, we know that at least one column of a Schur stable ${(A+BK)}$ has sum strictly smaller than 1. Further, any policy $K$ that results in a stable system ${(A+BK)}$ will drive the state (excluding $v_{g}$) to the origin. As $v_{g}$ is absorbing, all probability density must eventually be concentrated in $v_{g}$.
@@
where we use ${E=I}$, and ${\hat{r}\in\mathbb{R}^{n}}$ is the vector containing the element of $r$ (or zero if $u_{i}=0$) corresponding to the action chosen for each state. Further, for some choice of actions ${a\in\mathcal{A}(v)}$ for all ${v\in\mathcal{V}^{\prime}}$, the resulting transition matrix ${(A+BK)}$ attains the optimal cost of (1). This must then be the optimal cost also for the resulting SSP, since the cost is identical, with the corresponding actions giving a feedback law that solves (1). $\blacksquare$
@@
[…]
@@
$\mathcal{T}(v,a_{1})$ =[ 0  0.6  0.4  0]absentsuperscriptdelimited-[]$=\left[\;0\;\;0.6\;\;0.4\;\;0\;\right]^{\top}$
$\mathcal{T}(v,a_{2})$ =[ 0.3  0  0.7  0]absentsuperscriptdelimited-[]$=\left[\;0.3\;\;0\;\;0.7\;\;0\;\right]^{\top}$
$\mathcal{T}(v,a_{3})$ =[ 0  0.1  0.4  0.5]absentsuperscriptdelimited-[]$=\left[\;0\;\;0.1\;\;0.4\;\;0.5\;\right]^{\top}$
@@
Taking $x_{0}$ as the vector where ${x_{v_{0}}=1}$ and zero in all other elements gives the initial state. Extract ${\mathcal{V}=\mathcal{V}^{\prime}\setminus{G}}$ as the set of states in (1). The converse direction then follows directly from reversing the construction in Section 3.1 and applying the proof of Theorem 2. The optimal controller for (1) may not be unique. However, the controller given by (3) is always a solution to the SSP, attaining the same expected cost. $\blacksquare$ To the best of our knowledge, the equivalence resulting from Theorems 2 and 3, although intuitive, has not previously been observed in the literature. Despite this, there is much to be gained from making this connection explicit. Firstly, the success of methods developed to efficiently solve problems within the class of SSP can be utilized for optimal control of large-scale systems with continuous state and action spaces. The full class of problems on the form (1), without Assumption 2, is broader than the class of SSP as given in Definition 3.1. However, as is exemplified by the heuristic search algorithm presented in Section 4, tools and algorithms developed for SSP can be applied to all problems on the form (1), as the key advantages of the solution to the Bellman equation still hold. Secondly, the analysis provided in the optimal control framework provides a clear view of the problem class for which an explicit solution to the Bellman equation is obtained. This novel framework can be used to find previously unexplored classes of systems that share the computational advantages of SSP. One such example is the class of systems on the form of (1) with the summation constraints on the input vectors $u_{i}$ replaced by arbitrary $p$-norms. This setting is explored in (Li & Rantzer, 2023), although for a more restricted problem class with decoupled constraints.

### v2

@@ Abstract
This work considers infinite-horizon optimal control of positive linear systems applied to the case of network routing problems. We demonstrate the equivalence between Stochastic Shortest Path (SSP) problems and optimal control of a certain class of linear systems. This is used to construct a heuristic search framework for linear positive systems inspired by existing methods for SSP. We propose a heuristics-based algorithm for efficiently finding local solutions to the analyzed class of optimal control problems with a given initial state and positive linear dynamics. By leveraging the bound on optimality in each state provided by the heuristics, we also derive a novel distributed algorithm for calculating local controllers within a specified performance bound, with a distributed condition for termination. More fundamentally, the results allow for analysis of the conditions for explicit solutions to the Bellman equation utilized by heuristic search methods.
@@ 1 Introduction
The work presented here demonstrates the equivalence of two problem classes that are the focus of optimal control and graph search algorithms respectively. The first is a class of optimal control problems involving positive linear systems with linear cost and homogeneous constraints. The second is the class of Stochastic Shortest Path (SSP) problems and their generalizations, which are the focus of contemporary work on heuristic search algorithms, see (Bonet and Hansen, 2010) for a summary. In bridging the gap between these two formulations, we show that many results derived for either can be readily applied to the other.
Specifically, we give a framework for a heuristics-based solution of optimal control problems with continuous state and action space that belong to the above described class of positive systems. In a similar spirit to heuristic search algorithms developed for SSP, these methods are computationally feasible for high-dimensional state and action spaces. Additionally, we show that value iteration using both upper and lower heuristic bounds can leverage this measure of suboptimality to decentralize calculation with guarantees on global performance. The analogy between the two classes also gives an analytic framework for further development of existing methods for SSP using results within the theory of optimal control.
@@
The method of dynamic programming, developed in (Bellman, 1957), shares a common origin with Dijkstra’s algorithm, see (Dijkstra, 1959). As emphasized in (Sniedovich, 2006), the line of argument in Dijkstra’s original work closely mirrors the optimality principle as first formulated by Bellman. Since their inception, these methods have both gone on to inspire a wide range of extensions, seeing great success in their respective fields of control theory and computer science. Although the lines of work within each discipline may seem disparate, they share the common feature of utilizing explicit solutions to the Bellman equation to achieve computational efficiency. It is this feature that enables the present work to give favorable results for positive systems and interpret these in terms of their counterpart in the SSP formulation.
@@
A natural extension to Dijkstra’s algorithm is proposed in (Hart, Nilsson, and Raphael, 1968). The authors introduce a heuristic function that lower bounds the cost in any state and uses this to guide the search, prioritizing directions that give a low heuristic cost. Already in (Bertsekas and Tsitsiklis, 1991) the theoretical results of dynamic programming obtained in the deterministic case are extended to stochastic problems. Algorithms that implement similar ideas beyond the deterministic case through the use of heuristics are the subject of subsequent work in the field, see the seminal paper (Barto, Bradtke, and Singh, 1995). The close underlying connection between the study of heuristic functions and Lyapunov analysis is highlighted in (Perkins and Barto, 2001). Comprehensive theoretical results on control of the broader class of constrained Markov Decision Processes (MDP), which includes problems like SSP, can be found in the monograph (Altman, 1999). Therein, arguments based in dynamic programming and Lyapunov theory are used to formulate linear programs for problems in this class under certain conditions.
@@
In (McMahan, Likhachev, and Gordon, 2005) the authors incorporate the use of a heuristic upper bound and provide methods to find and improve such a bound. The extension is motivated by computational advantages, since the total cost is known to be upper bounded from the problem formulation of SSP. In this paper, the upper bound will additionally serve the purpose of guaranteeing the stability of solutions, which is not otherwise certain in the setting of optimal control. Experimental results demonstrating the robust performance of heuristic search methods for problems that deviate from the prescribed structure can be found in (McMahan and Gordon, 2005), together with an extended discussion of priority calculations for this family of algorithms. (Smith and Simmons, 2006) seek to improve existing algorithms by using upper and lower bound heuristics to inform the search. The class of problems for which heuristic methods can be applied is expanded in (Kolobov, Mausam, Weld, and Geffner, 2011) by the introduction of a broader class of SSP and a corresponding algorithm. The work (Schmoll and Schubert, 2018) builds on these previous results, further improving existing methods for certain classes of problems that can be rewritten as SSP.
@@
Within the field of optimal control, recent works like (Blanchini, Bolzern, Colaneri, De Nicolao, and Giordano, 2023) and (Rantzer, 2022) focus on explicit solutions to the Bellman equation for subclasses of positive linear systems. The latter work formulates a linear program for the synthesis of the optimal controller. Previously, it has been shown in (Rami and Tadeo, 2007) that stabilizing controller gains for linear positive systems can be found via linear programming. In (Ohlin, Tegling, and Rantzer, 2024), the class of applicable systems is extended to include coupled homogeneous input constraints, allowing for the modeling of network routing problems while maintaining these computational advantages. Similar properties of solutions to the Bellman equation for a set of MDP with cost function based on the Kullback-Leibler divergence are noted in (Todorov, 2006), with applications to optimal transport on networks. The use of upper and lower bounds on a general cost function for optimal control problems is explored in (Lincoln and Rantzer, 2006) as a means to relax the problem formulation, allowing solutions to be suboptimal to an extent, in exchange for improved performance.
@@ 1.1 Notation
Inequalities are applied element-wise for matrices and vectors throughout. When applied to a set, the notation $|\cdot|$} is taken to mean the cardinality of that set. Further, the notation $\mathbb{R}^{n}_{+}$ is used to denote the closed positive orthant of dimension $n$. The operator $\min\{{v},0\}$ extracts the minimum element of a vector $v$, yielding zero if v{v} has no negative elements. Let $A^{-\top}$ denote the inverse transpose of the matrix $A$. The expressions $\mathbf{1}_{p\times q}$ and $\mathbf{0}_{p\times q}$ signify a matrix of ones or zeros, respectively, of the indicated dimension, with subscript omitted when the size is clear from the context. If the dimension is zero, this is to be interpreted as the empty matrix. The matrix $A$ is said to be row or column substochastic if it holds that $A\mathbf{1}\leq\mathbf{1}$} or $\mathbf{1}^{\top}A\leq\mathbf{1}^{\top}$} respectively.
@@ 2 Problem setup
with $x_{0}\in\mathbb{R}^{n}_{+}$}. The input signal $u\in\mathbb{R}^{m}$} is partitioned into $n$ subvectors $u_{i}{\in\mathbb{R}^{m_{i}}}$ (possibly $m_{i}=0$}, implying that the corresponding $B_{i}$ is empty), so that $m=\sum_{i=1}^{n}m_{i}$}. The matrix $B=\begin{bmatrix}B_{1}\cdots B_{n}\end{bmatrix}\in\mathbb{R}^{n\times m}$} and $A\in\mathbb{R}^{n\times n}$} with $B_{i}\in\mathbb{R}^{n\times m_{i}}$} define the linear dynamics, with each submatrix $B_{i}$ determining the effect of the inputs $u_{i}$ associated with the $i$th constraint. Let the matrices $B_{i}$ be further subdivided into individual columns as $B_{i}=\begin{bmatrix}B_{i1}\cdots B_{im_{i}}\end{bmatrix}$} with $B_{ij}\in\mathbb{R}^{n}$}. The costs connected to the states and actions are $s\in\mathbb{R}^{n}_{+}$} with s>0{s>0} and $r\in\mathbb{R}^{m}_{+}$} with $r_{i}\in\mathbb{R}^{m_{i}}_{+}$} following the partition of $u$ and $r_{ij}$ denoting the $j$th element of $r_{i}$. The constraints on the input signal $u(t)$ are given by the matrix $E=\begin{bmatrix}E_{1}\cdots E_{n}\end{bmatrix}^{\top}\in\mathbb{R}^{n\times n}_{+}$}. The special case of singular $E$ typically corresponds to degenerate formulations and will not be treated here. Define the set $\mathcal{V}$ to contain the indices $\{1,\ldots,n\}$. Let $K=\left[K_{1}^{\top}\cdots K_{n}^{\top}\right]^{\top}$} be a feedback matrix with $K_{i}\in\mathbb{R}^{m_{i}\times n}_{+}$} and define
@@
The above formulation with strictly positive $s$ ensures that the system is observable (only $x=0$ yields zero immediate cost) with the specified cost function, meaning that the optimal feedback law must also be stabilizing if the total cost is finite. A more intricate formulation of this criterion is given in (Li and Rantzer, 2023). In the present setting, however, any observable positive system on the form (1) can be transformed to an equivalent problem satisfying $s>0$ by reallocation of the costs $s$ and $r$.
Let the function $J^{*}(x)$ give the optimal cost of an optimization problem initialized at $x$. For a general problem with immediate cost $g(x(t),u(t))$ and dynamics x(t+1)=f(x(t),u(t)){x(t+1)=f(x(t),u(t))}, the celebrated Bellman equation
$J^{*}(x)=\min_{u}\left[g(x,u)+J^{*}(f(x,u))\right]$} (2)
gives a condition to validate the optimality of $J^{*}$. The main tool used to find optimal solutions in the present proofs is the method of value iteration, which corresponds to fixed point iteration of (2). Of particular use is the property that value iteration converges monotonically to the fixed point of (2) under light assumptions on the initial value. For further reading, see (Bellman, 1957).
@@ 2.1 Positive networks with linear cost
for all $K\in\mathcal{K}$ and all states $x\in\mathbb{R}^{n}_{+}$.
@@
(Ohlin et al., 2024) Under Assumption 1, the following three statements are equivalent:
  • (ii) The problem (1) has a finite value for $x_{0}\in\mathbb{R}^{n}_{+}$.
  • (iiii) There exists $p\in\mathbb{R}^{n}_{+}$ satisfying the equation
@@
  • (iiiiii) The value of the linear program
@@
$p=s+A^{\top}p+\min_{K\in\mathcal{K}}\left[K^{\top}(r+B^{\top}p)\right]$ (5)
where the dependence on the constraint matrices $E_{i}$ is built into the definition of the set $\mathcal{K}$. The optimal cost is attained by choosing $K$ as in (4). Although less explicit, this expression is more compact and will be used when convenient in the treatment of the heuristic search algorithm presented in Sections 4 and 5 to illustrate the use of different policies for subsets of the state space $\mathcal{V}$.
@@ 3 The equivalent problem
The extension of previous results on positive linear systems to the class of systems covered by Theorem 1 lets us make a formal connection to other existing frameworks for network routing problems. As the definitions used in the literature concerning SSP (e.g. (Bertsekas and Tsitsiklis, 1991), (McMahan et al., 2005) and (Kolobov et al., 2011)) differ in interpretation, the following definition strives to be as broad as possible while maintaining applicability of the results in Theorem 1.
@@ A Stochastic Shortest Path (SSP) problem is defined by the following properties:
    a cost function $\mathcal{C}(v,a)$ associating each action with a nonnegative cost that satisfies $\mathcal{C}(v,a)=0$ for $v\in\mathcal{G}$ and $\mathcal{C}(v,a)>0$ for $v\in\mathcal{V}^{\prime}\setminus\mathcal{G}$
@@
Note that the requirement $\mathcal{C}(v,a)>0$ for $v\in\mathcal{V}^{\prime}\setminus\mathcal{G}$ is equivalent to assuming observability in the setting (1). (Kolobov et al., 2011) lifts this restriction to allow zero-cost cycles in states outside $\mathcal{G}$ to be part of the solution. Such cases require special treatment and will not be considered in the present work.
@@ 3.1 Reformulation
The problem (1) contains the class of SSP. In order to show this, we first introduce a helpful lemma, as well as an assumption on the system dynamics. The following lemma simplifies the subsequent problem construction by allowing the choice $E=I$ when the original $E$ is invertible.
@@
Consider a similarity transform $\hat{x}=Tx$ of the original state vector in (1). This gives rise to transformed constraints $^{\mathbf{1}^{\top}u_{i}\leq E_{i}^{\top}T^{-1}\hat{x}}$. Enforcing $ET^{-1}=I$} yields T=E{T=E}, which is a valid choice since $E$ is invertible. The original state $x$ is confined to $\mathbb{R}^{n}_{+}$ under the dynamics of (1). Further, the matrix $E$ is nonnegative by construction. Hence, it holds that $\hat{x}\geq 0$} starting from $\hat{x}_{0}=Tx_{0}$}. $\blacksquare$
Next, we introduce an assumption on substochasticity of the closed loop system in (1). As is shown below, both classes can be interpreted as SSP.
@@
The matrix (A+BK){(A+BK)} is column substochastic for all $K\in\mathcal{K}$.
Without this assumption, the problem can only be reformulated as an equivalent SSP under a relaxation of Definition 3.1, to allow for an infinite extension of the state space, as detailed below in Section 3.3. The resulting problem can, however, still be formulated as an optimization over finitely many actions, and the search methods introduced in subsequent sections, inspired by methods for SSP, are equally applicable regardless of this condition.
@@
[…]
@@ 3.2 Equivalence
Consider a problem on the form (1) with finite value in which the system satisfies Assumptions 1 and 2. Let $E$ be invertible. The optimal control problem can be rewritten as an SSP according to Definition 3.1, with equal expected cost. The optimal solution to this SSP is also an optimal control law for the problem (1).
@@
By Lemma 3.2, it is sufficient to consider systems with $E=I$. Properties 1-6 of Definition 3.1 are directly fulfilled by the construction in Section 3.1. All that remains is to show that property 7 also holds for the resulting SSP, that the expected cost is identical to that of the original problem and that the solution attains the optimal value of (1).
In order to show property 7 of Definition 3.1 we apply the method used in Chapter 2.3 of (Bertsekas and Tsitsiklis, 1996) to rewrite any discounted MDP as an SSP. In the system x(t+1)=(A+BK)x(t){x(t+1)=(A+BK)x(t)}, the quantity $\mathcal{T}(v,a)||_{1}$ corresponds to the column sums of (A+BK){(A+BK)}. The matrix (A+BK){(A+BK)} is Schur stable for some choices of $K\in\mathcal{K}$}, since the problem (1) has a finite value. In particular, this is at least true for the optimal feedback law. By the Perron-Frobenius theorem, we know that at least one column of a Schur stable (A+BK){(A+BK)} has sum strictly smaller than 1. Further, any policy $K$ that results in a stable system (A+BK){(A+BK)} will drive the state (excluding $v_{g}$) to the origin. As $v_{g}$ is absorbing, all probability density must eventually be concentrated in $v_{g}$.
@@
where we use E=I{E=I}, and $\hat{r}\in\mathbb{R}^{n}$} is the vector containing the element of $r$ (or zero if $u_{i}=0$) corresponding to the action chosen for each state. Further, for some choice of actions $a\in\mathcal{A}(v)$} for all $v\in\mathcal{V}^{\prime}$}, the resulting transition matrix (A+BK){(A+BK)} attains the optimal cost of (1). This must then be the optimal cost also for the resulting SSP, since the cost is identical, with the corresponding actions giving a feedback law that solves (1). $\blacksquare$
@@
Consider a problem on the form (1) with $n=3$, $m=4$, $x_{0}=\left[2\;0\;1\right]^{\top}$ and dynamics
@@
Constraints on the form (1) with $m_{1}=1$}, $m_{2}=2$}, $m_{3}=1$} and E=I{E=I} ensure positivity of the dynamics, fulfilling Assumption 1. Following the construction in Section 3.1, we get a state space $\mathcal{V}^{\prime}$ such that $|\mathcal{V}^{\prime}|=4$} where the fourth state acts as fictive goal state. Let $v$ be the state that corresponds to $x_{2}$. We then have $|\mathcal{A}(v)|=3$} with actions $a_{1}$, $a_{2}$ and $a_{3}$ corresponding to $u_{2}=\left[\;0\;0\;\right]^{\top}$}, $u_{2}=\left[E_{2}^{\top}x\;\;0\right]^{\top}$} and $u_{2}=\left[0\;\;E_{2}^{\top}x\right]^{\top}$} (the vertices in Figure 1). The cost function associated with these actions has value $\mathcal{C}(v,a_{1})=s_{2}=1$}, $\mathcal{C}(v,a_{2})=s_{2}+r_{21}=2$} and $\mathcal{C}(v,a_{3})=s_{2}+r_{22}=2$}. The transition function is constructed from the second column of $(A+BK_{a})$} for $a\in\mathcal{A}(v)$}. Should the transition probabilities sum to less than one, the remainder is allocated to the transition from $v$ to the goal state. This gives
@@
[…]
$\mathcal{T}(v,a_{3})$ =[ $=\left[\;0\;\;0.1\;\;0.4\;\;0.5\;\right]^{\top}$
@@
Taking $x_{0}$ as the vector where $x_{v_{0}}=1$} and zero in all other elements gives the initial state. Extract $\mathcal{V}=\mathcal{V}^{\prime}\setminus{G}$} as the set of states in (1). The converse direction then follows directly from reversing the construction in Section 3.1 and applying the proof of Theorem 2. The optimal controller for (1) may not be unique. However, the controller given by (4) is always a solution to the SSP, attaining the same expected cost. $\blacksquare$
To the best of our knowledge, the equivalence resulting from Theorems 2 and 3, although intuitive, has not previously been observed in the literature. Despite this, there is much to be gained from making this connection explicit. Firstly, the success of methods developed to efficiently solve problems within the class of SSP can be utilized for optimal control of large-scale systems with continuous state and action spaces. The full class of problems on the form (1), i.e. without Assumption 2, is broader than the class of SSP as given in Definition 3.1. However, as is exemplified by the heuristic search algorithm presented in Section 4, tools and algorithms developed for SSP can be applied to all problems on the form (1), as the key advantages of the solution to the Bellman equation still hold. Secondly, the analysis provided in the optimal control framework provides a clear view of the problem class for which an explicit solution to the Bellman equation is obtained. This novel framework can be used to find previously unexplored classes of systems that share the computational advantages of SSP. One such example is the class of systems on the form of (1) with the summation constraints on the input vectors $u_{i}$ replaced by arbitrary $p$-norms. This setting is explored in (Li and Rantzer, 2023), although for a more restricted problem class with decoupled constraints.